A galvanometer has an internal resistance of 100 Ω and deflects full-scale at 2.00 mA. What size resistor should be added to it to convert it to a milliammeter capable of reading up to 4.00 mA? Please explain the answer.
Answer #1

R1 = 100 ohm

I1 = 2.00 mA

Which gives V = 2*10^-3*100 = 0.2 V

Since V will be same

Req = V/I = 0.2/0.004 = 50 ohm

So we need to 100 ohm resistor in parallel with the given resistor so that Req can be 50 ohm

Answer will be 100 ohm in parallel